OpenStudy (anonymous):

evaluate the double integral: xcos(y)DA, D is bounded by y=0, y=x^2, x=1

OpenStudy (anonymous):

\[\int\limits \int\limits xcos(y)\]

OpenStudy (anonymous):

i got to \[\int\limits_{0}^{1}xsin(x^{2})dX\]

OpenStudy (anonymous):

then I'm stuck

OpenStudy (turingtest):

u-sub\[u=x^2\implies du=2xdx\]

OpenStudy (anonymous):

so 1/2(cos x^2) from 0 to 1?

OpenStudy (anonymous):

figured it out! thanks!

OpenStudy (turingtest):

-1/2cos, but yeah

OpenStudy (mendicant_bias):

So does this just yield \[\frac {1}{2} - \frac {1}{2}\cos(1)\] ?